

********************************************************************
UARS WEEKLY STATUS REPORT - 10 APRIL 1998
********************************************************************
***********
SOLSTICE
***********
SOLSTICE continues to operate normally, collecting solar data on all
available orbital tracking periods and stellar calibration observations,
whenever possible.
***********
SUSIM
***********
SUSIM successfully gathered its normal solar, offset, and occultation
scan data this week without apparent problems. SUSIM has recently begun
taking an second scan of the wavelength region containing the Mg II
absorption feature. It is expected that the noise level of the Mg II
index will be reduced though the inclusion (at some time in the future)
of this new data in its calculation.
***********
PEM
***********
PEM is currently off.
PEM is awaiting evaluation by GSFC of its proposed activation procedure.
The minimal configuration for PEM AUX activation is the zenith VMAG, MEPS1,
MEPS2, and HEPS1 units, only. This configuration is driven by the capabilities
of the spacecraft OBC and may increase if OBC program area increases. PEM
activation is proposed to be OBC driven when there exists plenty of power to
operate PEM. It is anticipated that the OBC will activate PEM on the daylight
portions of every orbit.
***********
HALOE
***********
HALOE was powered on April 8, 1998. All engineering parameters are nominal
and the science channels are within the expected operating range.
***********
MLS
***********
The MLS instrument has been powered off since Mar 29. MLS is scheduled to
begin north viewing operations on Apr 15.
***********
HRDI
***********
HRDI is currently operating. We expect to be operating till April 15.
Operations are normal.
***********
WINDII
***********
Nothing received.
***********
MPG
***********
Routine processing of instrument Daily Activity Plans continues
with no problems. The MPG continues to assist the Project
Scientist with the scheduling of instrument operations.
***********
CDHF
***********
The CDHF has received a total of 381 CD-ROM requests as of 10-April-1998.
There were no orders received this week.
The CDHF processed current day data through level 3 for all active
instruments. Additionally, CDHF personnel reprocessed: (R)= Requested
(C) = Complete, (F)= Failed (L) = Days remaining to reprocess.
25 days PEM (3.32) PROD (1379R-1099C=280L)
54 days MLS (4.90) PROD (1297R-1279C-18F=0L)
5 days CLAES (7.0) ENG (72R-40C-1F=31L)
2178 days ATT (7.2) PROD (2178R-439C=1739L)
The CDHF completed setup of database compression, to run during GRAVE shifts.
This moves the UCSS downtime for partial compresses to offshift hours, causing
less disruption to daytime users for UCSS downtime. Full compresses will
still be performed once a month on a scheduled Tuesday. Archives with UCSS
up, are still performed on Thursdays.
The CDHF developed software to assist with the monitoring of the UARS RACs.
When a node has been down for an hour, CDHF operators will send e-mail to a
distribution list, requesting that the NISN/NI NOC open a trouble ticket and
investigate.
The CDHF continued work on network and RAC configurations. The reports
include contacts, decnet addresses, ip addresses, locations, phone numbers,
and email adresses. Work also continues in providing NISN the RACs that
need Multinet, and writing up the UARS Network Requirements.
***********
FOT
***********
REPORTING PERIOD
This report covers 28 March 1998 (Orbit 35750, GMT Day 087) through 03
April 1998 (Orbit 35853, GMT Day 093).
SPACECRAFT OPERATION
The observatory is now in Reverse Flight (Northern Hemisphere viewing) and
is performing nominally under two-battery operations. The instrument
operational changes were:
INSTRUMENT OPERATIONAL CHANGES
Instrument Time Orbit Comment
MLS 089/0136 35780 Off
The beta angle reached a minimum angle of 0.0 degrees on 02 April 1998 (DOY
092). The beta angle is now increasing toward a maximum angle of 45.7
degrees on 19 April 1998 (DOY 109).
A Forward-to-Reverse Yaw Around maneuver (number 67) was successfully
executed in Orbit 35833 at 13:47:22 UTC on 02 April 1998. The SSPP
Instruments (ACRIM, SOLSTICE, and SUSIM) were enabled. Post maneuver
analysis of data indicate that during the Yaw Around, the end-of-night load
bus voltage was 24.5 volts and the average depths of discharge (DODs) on
Batteries 2 and 3 were approximately 22.5 and 22.0 percent respectively.
Spacecraft battery 1, 2, and 3 performance monitoring continues. Battery 1
remains OFF the charge relay. Battery 2 and Battery 3 are currently
maintaining greater than 24.5 V end-of-night (EON) load bus voltage.
Battery temperatures are stable with a temperature delta between Battery 2
and Battery 3 of 1.1 degrees C. Additional battery performance data for
each battery on 28 March 1998 and 03 April 1998 are provided in Appendix D.
The clock error ranged from -12.8 to +10.0 msec during this report period.
Clock rate adjustments are listed in Appendix C.
All instrument and subsystem engineers were advised on the initial
excursion of ALERTS or Out-Of-Limits conditions detected by the FOT for
this report. A summary is listed below and the details of ALERT and Out-
Of-Limits occurrences are listed in Appendix B.
% ALERTS = none
% OUT-OF-LIMITS = 08
The following real-time command sheets were executed for the listed
operational element(s) during this report period (see Appendix A).
% FOT 02
% HRDI 02
% MLS 01
% SUSIM 01
UARS Anomaly Reports initiated during this report period are listed in
Appendix E and summarized below.
% FOT 01
One (1) Anomaly Reports remained OPEN at the end of this report period (see
Appendix E).
TRANSPONDER FREQUENCIES:
Transponder A: 98/062 - 2287.496945
Transponder B: 98/062 - 2287.497010
GROUND SYSTEM OPERATION
One (1) operational support problem(s) occurred this period. Anomaly
Reports are included in Appendix E.
The data loss calculations supplied by Data Capture Facility (DCF) from
mission start through UTC Day 98/085 show the observatory data loss to be
20 hours, 42 minutes, 33.140 seconds (no increase since last report
period). This is a 0.0362 percent data loss which equals a 99.9638 percent
data capture for the mission.
The FOT attempted to perform return link IP testing this week, but was
unable to do so due to configuration problems with the UARS/MSOCC IP
conversion device. Further testing will be attempted next week. In
addition, the FOT is in the process of finalizing the configuration
requirements for the IP Forward link test which will utilize the UTTS
instead of the actual UARS spacecraft. To accommodate this test, the
network will have to be configured in a command loop-back mode so commands
sent by the FOT will travel through the IP network, up to TDRS, through the
RFSOC, and to the simulator. Actual testing of this configuration is
expected to occur within the next two weeks. A deadline for complete
switch-over to the IP network has been set for mid July 1998.
The FOT continued making modifications to the OBCSTF resident POCC
database. These changes will be incorporated into the next operational
database release, version 66.
OTHER SIGNIFICANT EVENTS
Work on the upgrade of the OBC Power Monitor (PMON) software continued.
Continued database work for new PMON software. Defined new RTS and PDB
assignments to support PEM day-only operation. Five new RTS allocations
will be required, holding 179 commands. Three previously allocated, but
unused, RTSs, and eight unused PDBs will also be used. These new RTSs and
PDBs will be included with the new PMON software uplink.
FOT continues to support CNMOS in their efforts to make necessary software
changes (MSOCC, CMS, DCF, etc.) to support the year 2000 rollover.
Monitoring of the magnetic field model performance continues.
FOT continues to improve a Power Profile for UARS based on the current
status of the Modular Power System.
FOT continues to support the new UARS ground system re-engineering effort.
MEETINGS
No UARS Battery telecon was held this week. The next telecon has not been
scheduled.
FUTURE SIGNIFICANT EVENTS
The next UARS Yaw Around Maneuver is currently scheduled for 05 May 1998.
A second building 23 power outage is scheduled from about 12 p.m., 24 April
1998, until about 12 p.m., 27 April 1998.
UARS Flight Operations plans to attend the Program Experience portion of
the Space Power Workshop at JPL in April 1998. UARS FOT plans on
presenting an overview of the UARS power subsystem and its performance.
The UARS orbit altitude is dropping slowly and a drag makeup orbit adjust
is targeted for 05 August 1998.
PERSONNEL CHANGES
None.
Attachments: Appendix A, Real-time Commands Sheets
Appendix B, Out-Of-Limits & ALERT Occurrences
Appendix C, Clock & Solar Array Rate Corrections
Appendix D, UARS Battery Performance Data
Appendix E, UARS Anomaly Reports
APPENDIX A
REAL-TIME COMMAND SHEETS
Week ending 03 April 1998
ORBIT AOS NO. SUBSYS REASON
----- --- --- ------ -------
35778 088/2233 98-117 OBC Clk Adj fr 0.00 to 20.6 msec/day
35780 089/0136 98-118 MLS Power OFF
35796 090/0321 98-119 OBC Clk Adj fr 20.6 to 0.00 msec/day
35801 090/1106 98-120 HRDI HRDI Heater Delay to 125 Amp/Min
35813 091/0513 98-121 HRDI HRDI Heater Delay to 175 Amp/Min
35839 092/2328 98-122 SUSIM reload instrument commands
_______________________________________________________________
APPENDIX B
OUT-OF-LIMITS & ALERTS OCCURRENCES
Week ending 03 April 1998
MNEMONIC STATE DESCRIPTION ORBITS
--------- ----- ------------ ------
HRIFESURVTMP Y-HI IFE Survival Heater Temp 35750-1
HRRADTEMP Y-HI IFE Radiator Temp 35751
QTSISAIF Y-HI SSPP Baseplate SISA I/F 35756-853
HRPSEQBIMON Y-HI PSE Quiet Bus Current Monitor 35771
SOSTMPRNTRN Y-LO Solstice Entrance Temperature 35821,31,36
ACTEMPANA2 Y-LO Temp Sensor Electronics 35831-36
HRPSEQBIMON Y-HI PSE Quiet Bus Monit 35831
WIEUTEMP Y-LO WINDII EU Temperature 35750-853
The following ALERT occurrences were detected:
MNEMONIC DESCRIPTION ORBITS
--------- ------------ ------
None this report period
APPENDIX C
CLOCK & SOLAR ARRAY RATE CORRECTIONS
Week ending 03 April 1998
CLOCK DRIFT RATE ADJUSTMENTS
ORBIT TIME CLOCK RATE CHANGE
----- ---- -----------------
35778 088/22:38 (29 Mar 98) from 00.0 to 20.6 msec/day
35796 090/03:26 (31 Mar 98) from 20.6 to 00.0 msec/day
_______________________________________________________________
APPENDIX D
UARS BATTERY PERFORMANCE DATA
Week ending 03 April 1998
28 Mar 1998 - GMT Day 087 Beta = 18.9 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-4.31 +0.0/ +0.0 18.4 0.000 0.0
2 +84.0/ +0.0 +7.09/+5.47 +31.6/-14.4 24.5 1.025 23.1
3 +67.2/ -5.5 +5.47/+4.19 +30.4/-16.0 24.5 1.024 22.5
03 Apr 1998 - GMT Day 093 Beta = 4.5 deg, SA Pos = 269 deg
DIFF V (mV) TEMP (C) CURR (amp) EON V AVE C/D AVE
BAT MAX/MIN MAX/MIN MAX/MIN MIN RATIO DOD(%)
--- ------------ ---------- ----------- ------- ------- ------
1 +728.0/+728.0 -3.05/-3.99 +0.0/ +0.0 18.4 0.000 0.0
2 +112.0/ +0.0 +5.79/+4.51 +31.2/-14.4 24.5 1.024 22.4
3 +84.0/ -5.6 +4.51/+3.55 +30.4/-15.2 24.5 1.021 21.8
V/T LEVEL CHANGES
ORB# TIME LVL FR-TO
----- -------- ----------
no changes
APPENDIX E
UARS ANOMALY REPORTS
Week ending 03 April 1998
NEW ANOMALY REPORTS GENERATED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-014 FOT 35830 092/08:48 Tac 4 Crash
ANOMALY CLOSURE INFORMATION RECEIVED
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-009 FOT March 1998 14170 Genric Late Acq.
98-014 FOT 35830 092/08:48 Tac 4 Crash
ANOMALY INVESTIGATION REPORTS GENERATED
AIR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------- ------ ----- ---- --- -----------------
None this report period
OPEN ANOMALY REPORTS
AR NO. SUBSYS ORBIT TIME TTR PROBLEM/DATA LOSS
------ ------ ----- ---- --- -----------------
98-015 Generic Late Acq for April 1998. Open until 31 April 1998.
SEVNTFW13